In the linear system 2m+5n=14 2m-3n=6, how do I solve using the substitution method? Which variable do you recommend I begin with?
Answer by stanbon(75887)

In the linear system 2m+5n=14 2m-3n=6, how do I solve using the substitution method? Which variable do you recommend I begin with?
-----------------------------------------------
2m+5n=14
2m-3n=6
-----
It's your option.
----
Solving for "m" you get:
m = (14-5n)/2
---
Substitute for "m" and solve for "n":
2[(14-5n)/2] - 3n = 6
------
14-5n - 3n = 6
8n = 8
n = 1
---------
Solve for "m":
2m - 3n = 6
2m - 3 = 6
2m = 9
m = 9/2
================
Cheers,
Stan H.
================